OpenStudy (anonymous):

How do you find the center of a circle with a equation if you are given (x-4)^2+(y-2)^2=25?

OpenStudy (whpalmer4):

with great ease :-) formula of a circle with center \((h,k)\) and radius \(r\) is \[(x-h)^2+(y-k)^2=r^2\]Compare that with your equation:\[(x-4)^2+(y-2)^2=25\]